OpenStudy (yacoub1993):

Find the area of a triangle with a base of 8 m and a height of 6 m. A. 7 m2 B. 48 m2 C. 24 m2 D. 14 m2

OpenStudy (shubhamsrg):

What's your attempt ?

OpenStudy (yacoub1993):

\[24m ^{2}\]

OpenStudy (anonymous):

Area of triangle= 1/2 * base * height

OpenStudy (anonymous):

You are correct.

OpenStudy (yacoub1993):

8*6=48/2 was my attempt

OpenStudy (yacoub1993):

Another Question

OpenStudy (shubhamsrg):

area of trapezium = 1/2 (sum of parallel bases)*(height)

OpenStudy (yacoub1993):

i have no idea of what to do here

OpenStudy (anonymous):

What @shubhamsrg said. :) that should help.

OpenStudy (yacoub1993):

is it 219.96

OpenStudy (yacoub1993):

\[219.96 yd ^{2}\]

OpenStudy (yacoub1993):

Another Question

OpenStudy (yacoub1993):

i think it should be \[67.94m ^{2}\]

OpenStudy (yacoub1993):

is my answer correct guys